OpenStudy (anonymous):

Looking for insight onto solving this integral. I know it requires some trig sub and and completing the square but I cant fig it out! thanks

OpenStudy (anonymous):

\[\int\limits_{}^{}\sqrt{8x-x^2}dx\]

zepdrix (zepdrix):

Hi c:

OpenStudy (anonymous):

Hey :D

zepdrix (zepdrix):

Completing the square? Hmmm. Personally I would start by taking the negative off of the square term. I think it'll just be easier to work with.\[\Large\bf\sf \int\limits \sqrt{-(x^2-8x)}\;dx\]And then let's complete the square on stuff in the brackets. That seem ok? :o

OpenStudy (anonymous):

Yea I tried that I get something like sqrt(16-(x-4)^2)

zepdrix (zepdrix):

Ah yes, very good.

zepdrix (zepdrix):

We would like our thing under the root to resemble something like this:\[\Large\bf\sf 1-(stuff)^2\]Because then we can make a nice substitution,\[\Large\bf\sf stuff=\sin \theta\]And apply our square identity to simplify it down,\[\Large\bf\sf 1-(stuff)^2\quad=\quad 1-\sin^2\theta\quad=\quad \cos^2 \theta\]

zepdrix (zepdrix):

But see how we have a 16 as our first value under the root? Uh oh it's not a 1. What can we do? :O

OpenStudy (anonymous):

Factor out a four \[\int\limits_{}^{} and thats where im not sure how to proceed

zepdrix (zepdrix):

We have to factor a 16 out of the square term as well, I'm going to bring it into the square as a 4, \[\Large\bf\sf 4\int\limits \sqrt{1-\left(\frac{x-2}{4}\right)^2}\;dx\]Does that make sense?

zepdrix (zepdrix):

x-4.. woops lemme fix that real quick.

zepdrix (zepdrix):

\[\Large\bf\sf 4\int\limits\limits \sqrt{1-\left(\frac{x-4}{4}\right)^2}\;dx\]

OpenStudy (anonymous):

OKay and then would I let x=sin?

zepdrix (zepdrix):

You let the big blob inside of the square = sin theta.

zepdrix (zepdrix):

\[\Large\bf\sf \frac{x-4}{4}\quad=\quad \sin \theta\]

zepdrix (zepdrix):

\[\Large\bf\sf 4\int\limits\limits\limits \sqrt{1-\left(\sin \theta\right)^2}\;dx\]Which cleans up the stuff under the root quite nicely, yes?

OpenStudy (anonymous):

Yea that does seem to work, and than I assume you can write \[4\int\limits_{}^{}\cos \theta d \theta \]

OpenStudy (anonymous):

=\[4\sin \theta \]

zepdrix (zepdrix):

No, our differential doesn't simply switch from dx to dtheta. He's going to give us some extra bagage with him.

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

so i find \[d \theta \] by taking the deravtive of our x=4sin(theta)+4

OpenStudy (anonymous):

dx=4cos(theta)d(theta)

zepdrix (zepdrix):

Ok looks good.

OpenStudy (anonymous):

\[16\int\limits_{}^{}(\cos \theta)^2d \theta \] ?

zepdrix (zepdrix):

Ok good.

zepdrix (zepdrix):

Remember your Cosine Half-Angle Identity?

OpenStudy (anonymous):

yea i believe the integral comes to, (theta/2)+sin(2theata)/4

zepdrix (zepdrix):

Ok good, with a 16 multiplying both terms yes?

OpenStudy (anonymous):

Yea correct

zepdrix (zepdrix):

So now we need our solution back in terms of x, which will take a little bit of effort.

zepdrix (zepdrix):

Theta doesn't have a nice x representation so for our first term we'll just attack it head on using our substitution from earlier.\[\Large\bf\sf \frac{x-4}{4}=\sin \theta\qquad\to\qquad \theta=\arcsin\frac{x-4}{4}\]

zepdrix (zepdrix):

\[\Large\bf\sf 8\theta+4\sin2\theta+\mathcal C\] So that will deal with the first term. Understand how I came up with the inverse sine?

OpenStudy (anonymous):

Yea because it is basically saying where does sin =

OpenStudy (anonymous):

But I find it confusing how to go about making the triangle

zepdrix (zepdrix):

Before we make a triangle, we need our sin2theta to be in terms of the angle theta, not 2theta. Remember Sine Double Angle? :o

OpenStudy (anonymous):

Sin2theta=2sincos

zepdrix (zepdrix):

Ok good,\[\Large\bf\sf 8\theta+8\sin\theta\cos\theta+\mathcal C\]

zepdrix (zepdrix):

Looking back at our substitution,\[\Large\bf\sf \sin \theta\quad=\quad\frac{\color{royalblue}{x-4}}{4}\quad=\quad \frac{\color{royalblue}{opposite}}{hypotenuse}\]There is one way we can set up our triangle.

zepdrix (zepdrix):



zepdrix (zepdrix):

Understand how to label the sides?

OpenStudy (anonymous):

Yup! but which sqrt do we use for the adjacent?

zepdrix (zepdrix):

Use Pythagorean Theorem to find the adjacent side.

zepdrix (zepdrix):

\[\Large\bf\sf b^2=4^2-(x-4)^2\]

OpenStudy (anonymous):

Alright makes sense, and after that just plugging in would the give the final result?

zepdrix (zepdrix):

Yes. For this problem you'll need to plug in both for your sin(theta) and your cos(theta). Shouldn't be too bad though.

OpenStudy (anonymous):

Yup, thanks for the help

zepdrix (zepdrix):

np

OpenStudy (anonymous):

And for the theta alone, what do i use? arcsin?

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

Yes, that's as good as we're going to get for our theta.\[\Large\bf\sf 8\arcsin\left(\frac{x-4}{4}\right)+8\sin\theta\cos\theta+\mathcal C\]

OpenStudy (anonymous):

Yea so 8arcsin(x-4/4)+8(x-4/4)(8-x/4)?

zepdrix (zepdrix):